OpenStudy (anonymous):

Find the coordinates of the midpoint of line segment AB for A(-3, 6) and B(1, 7).

OpenStudy (anonymous):

Midpoint formula is: (x1+x2)/2, (y1+y2)/2 Plug in the numbers: (1+(-3))/2, (6+7)/2 Simplify: -2/2 = -1 13/2 = 13/2 or 6.5 So the midpoint is at (-1,6.5)

OpenStudy (anonymous):

Thank you so much!